# Large Language Models as In-context AI Generators for Quality-Diversity

### v1

@@ Abstract
Quality-Diversity (QD) approaches are a promising direction to develop open-ended processes as they can discover archives of high-quality solutions across diverse niches. While already successful in many applications, QD approaches usually rely on combining only one or two solutions to generate new candidate solutions. As observed in open-ended processes such as technological evolution, wisely combining large diversity of these solutions could lead to more innovative solutions and potentially boost the productivity of QD search. In this work, we propose to exploit the pattern-matching capabilities of generative models to enable such efficient solution combinations. We introduce In-context QD, a framework of techniques that aim to elicit the in-context capabilities of pre-trained Large Language Models (LLMs) to generate interesting solutions using the QD archive as context. Applied to a series of common QD domains, In-context QD displays promising results compared to both QD baselines and similar strategies developed for single-objective optimization. Additionally, this result holds across multiple values of parameter sizes and archive population sizes, as well as across domains with distinct characteristics from BBO functions to policy search. Finally, we perform an extensive ablation that highlights the key prompt design considerations that encourage the generation of promising solutions for QD.
@@ Introduction
Quality-Diversity (QD) Pugh et al., (2016); Cully and Demiris, (2017); Chatzilygeroudis et al., (2021) algorithms are an effective way to collect and maintain a diversity of the best solutions seen and has been a step towards open-ended algorithms. Instead of just searching for a single optimal solution in conventional optimization, QD searches for a diversity of high-performing solutions. A QD approach is useful when one wants a diversity of good options and solutions at the end of search, but can also be implicitly effective in the conventional optimization setting by providing stepping stones to a more optimal solution (Gaier et al.,, 2019; Clune,, 2019), that otherwise might have been discarded. This has benefits across many applications such as robotics (Cully et al.,, 2015), procedural content generation (Gravina et al.,, 2019), engineering design (Gaier et al.,, 2018) and many more.
@@
Figure 1: Overview of In-context QD using LLMs as idea and solution generators. The LLM observes the current state of the archive which contains high-quality and diverse solution examples and uses this to generate new solutions. The prompt consists of context and query portions which are concatenated together.
@@
Variations and Optimizers. In MAP-Elites, the mutation can be done using any common mutation operator such as standard GA (Mouret and Clune,, 2015) or specific operators tailored for QD (Vassiliades and Mouret,, 2018). Previous work has also proposed to use gradient-informed mutation when available (Fontaine and Nikolaidis,, 2021; Nilsson and Cully,, 2021; Pierrot et al.,, 2022), or mutation based on evolution strategies (Colas et al.,, 2020; Fontaine et al.,, 2020; Flageat Lim and Cully,, 2023). In comparison, our work leverages the pattern completion abilities of LLM to generate the next solution given the large diversity of the archive of elites. While LLMs have also been used as mutation operators for their ability to tackle text-based and program-based QD domainsLehman et al., (2023); Bradley et al., (2023), we focus and demonstrate these capabilities on general QD BBO optimization problems, other than text or code and develop prompt construction strategies specifically for QD. Also close to our work, BOP-Elites (Kent and Branke,, 2020) and DDE-Elites (Gaier et al.,, 2020) both learn a representation of the current archive content and use this representation as a mutation operator to generate new relevant offspring. They rely on a Gaussian Process and a Variational Autoencoder model respectively. While our work also uses a representation of the archive content to generate relevant offspring, it relies on the in-context capabilities of pre-trained foundation models and does not require training a model online during the optimization. We show in the following that the LLM manages to generate relevant offspring without requiring any online specific training.
@@
Prior work using LLMs as optimizers has mainly focused on either (i) conventional optimization problems where the goal is to find a single solution instead of a QD problem (Yang et al.,, 2023; Mirchandani et al.,, 2023; Meyerson et al.,, 2023; Zhang et al.,, 2023), or (ii) focuses on the QD problem but uses the LLM in a way that does not explicitly account for both the fitness and feature in the prompt and has mainly been used in subjective text-based domains (Bradley et al.,, 2023; Samvelyan et al.,, 2024). There is preliminary work which uses LLMs for more generic evolutionary optimization (Liu et al.,, 2023) but considers the case of conventional optimization and requires very detailed task information and descriptions. Closest to our approach is concurrently released work which also studies prompting strategies but targeting recombination operators for Evolution Strategies (ES) (Lange et al.,, 2024). Instead, our work considers the general QD problem which contains the added dimensions of diversity. Additionally, our approach taps into the curated solutions in the archive of elites to build the context and studies in-context prompt construction approaches for the QD problem.
@@
In QD, the goal is to find a diversity of high-quality solutions. This provides two dimensions of optimization, fitness and novelty/diversity, compared to conventional single-objective optimization approaches which only consider fitness. A common approach when using LLMs as optimizers is to sort the sequence in an increasing order of fitness and provide a query of higher value, allowing the model to complete the sequence (Yang et al.,, 2023; Mirchandani et al.,, 2023). However, in QD, we are not concerned with just finding a solution that has a high fitness, we are also interested in covering the feature space with high-scoring solutions. Figure 1 shows an overview of our approach to this, In-context QD. The LLM takes as input the current state of the archive, and uses this as a prompt to generate new solutions. In the following, we explain how we represent the solutions and build a prompt that consists of the context and the query. This provides a general and modular framework for which to use LLMs as AI generators for Quality-Diversity.
@@
The solutions maintained in a QD archive lend themselves nicely to few-shot prompting of LLMs as the diversity and quality of solutions provide a good set of examples and context to refer to in order to generate better ideas and solutions. Ideally, the entire archive could be used as the context. However, given the context size limit of many open-source LLMs, we use a sample/subset of solutions of the archive as an approximation of the archive to construct the context. If provided with models with larger context or problems with lower archive sizes, the entire archive could be used as the context. Yet, despite this temporary limitation, we can already demonstrate the effectiveness of this proposed approach.
In building the context, there are two considerations: the content and the structure of the context. As explained above, the content of the context is a representation of the archive. We obtain this by uniformly sampling from the contents of the archive. An increasingly larger context of solutions sampled would give us an increasingly better representation of the archive culminating in the entire archive itself. Meyerson et al., (2023) suggest that few-shot prompting can be viewed as Estimation of Distribution Algorithms (EDAs). In our QD setting, by providing more solutions from the archive in the context using our prompt template, we provide a good approximation of the patterns and the distributions of the fitness-feature space. This provides a better context for the language model especially when queried for desired features.
@@
The selection of features used in the query $d^{q}$ cannot be treated entirely the same as with the fitness. In the case of features, the goal is to cover the entire feature space and not find an optimal point in the space by moving in a specific direction of a feature. While expanding the horizon and boundary of the archive is needed, it is also just as important to fill in ”holes” in the archive to maximize the coverage of the feature space. An unbiased approach to this is to uniformly sample the centroids of the cells in the archive to obtain query features $d^{q}_{1},...,d^{q}_{G}$. We refer to this approach as ”Uniform”. We also propose a strategy which queries features from cells that are empty first after which we default to sampling features uniformly from the cells (i.e. centroids) in the archive when the number of empty cells becomes less than the batch size. We refer to this strategy as ”Empty”.
@@
We compare in-context QD against a Random Sampling baseline which samples the parameter space uniformly and MAP-Elites (Mouret and Clune,, 2015) with the iso-line variation operator (Vassiliades and Mouret,, 2018). For all our algorithms and baselines, we use a batch size of $10$ at each generation, and run for 1000 generations. We also investigate the performance across different feature space sizes $C\in\{400,1600\}$ by changing the resolution of the MAP-Elites archive $C\in\{20^{2},40^{2}\}$ to ensure the generality of our approach. To be able to run in-context QD on just a single GPU device, we use Mistral-7B-v0.1 (Jiang et al.,, 2023), an open-source 7B parameter language model.

### v2

@@ Abstract
Quality-Diversity (QD) approaches are a promising direction to develop open-ended processes as they can discover archives of high-quality solutions across diverse niches. While already successful in many applications, QD approaches usually rely on combining only one or two solutions to generate new candidate solutions. As observed in open-ended processes such as technological evolution, wisely combining large diversity of these solutions could lead to more innovative solutions and potentially boost the productivity of QD search. In this work, we propose to exploit the pattern-matching capabilities of generative models to enable such efficient solution combinations. We introduce In-context QD, a framework of techniques that aim to elicit the in-context capabilities of pre-trained Large Language Models (LLMs) to generate interesting solutions using few-shot and many-shot prompting with quality-diverse examples from the QD archive as context. Applied to a series of common QD domains, In-context QD displays promising results compared to both QD baselines and similar strategies developed for single-objective optimization. Additionally, this result holds across multiple values of parameter sizes and archive population sizes, as well as across domains with distinct characteristics from BBO functions to policy search. Finally, we perform an extensive ablation that highlights the key prompt design considerations that encourage the generation of promising solutions for QD.
@@ Introduction
Quality-Diversity (QD) (Pugh et al.,, 2016; Cully and Demiris,, 2017; Chatzilygeroudis et al.,, 2021) algorithms are an effective way to collect and maintain a diversity of the best solutions seen and has been a step towards open-ended algorithms. Instead of just searching for a single optimal solution in conventional optimization, QD searches for a diversity of high-performing solutions. A QD approach is useful when one wants a diversity of good options and solutions at the end of search, but can also be implicitly effective in the conventional optimization setting by providing stepping stones to a more optimal solution (Gaier et al.,, 2019; Clune,, 2019), that otherwise might have been discarded. This has benefits across many applications such as robotics (Cully et al.,, 2015), procedural content generation (Gravina et al.,, 2019), engineering design (Gaier et al.,, 2018) and many more.
@@
Figure 1: Overview of In-context QD using LLMs as idea and solution generators. The LLM observes the current state of the archive which contains high-quality and diverse solution examples and uses this to generate new solutions to be evaluated and considered for addition to the archive. The prompt consists of context and query portions which are concatenated together.
@@
Variations and Optimizers. In MAP-Elites, the mutation can be done using any common mutation operator such as standard GA (Mouret and Clune,, 2015) or specific operators tailored for QD (Vassiliades and Mouret,, 2018). Previous work has also proposed to use gradient-informed mutation when available (Fontaine and Nikolaidis,, 2021; Nilsson and Cully,, 2021; Pierrot et al.,, 2022), or mutation based on evolution strategies (Colas et al.,, 2020; Fontaine et al.,, 2020; Flageat Lim and Cully,, 2023). In comparison, our work leverages the pattern completion abilities of LLM to generate the next solution given the large diversity of the archive of elites. While LLMs have also been used as mutation operators for their ability to tackle text-based and program-based QD domains(Lehman et al.,, 2023; Bradley et al.,, 2023), we focus and demonstrate these capabilities on general QD BBO optimization problems, other than text or code and develop prompt construction strategies specifically for QD. Also close to our work, BOP-Elites (Kent and Branke,, 2020) and DDE-Elites (Gaier et al.,, 2020) both learn a representation of the current archive content and use this representation as a mutation operator to generate new relevant offspring. They rely on a Gaussian Process and a Variational Autoencoder model respectively. While our work also uses a representation of the archive content to generate relevant offspring, it relies on the in-context capabilities of pre-trained foundation models and does not require training a model online during the optimization. We show in the following that the LLM manages to generate relevant offspring without requiring any online specific training.
@@
Prior work using LLMs as optimizers has mainly focused on either (i) conventional optimization problems where the goal is to find a single solution instead of a QD problem (Yang et al.,, 2023; Mirchandani et al.,, 2023; Meyerson et al.,, 2023; Zhang et al.,, 2023), or (ii) focuses on the QD problem but uses the LLM in a way that does not explicitly account for both the fitness and feature in the prompt and has mainly been used in subjective text-based domains (Bradley et al.,, 2023; Samvelyan et al.,, 2024). There is preliminary work which uses LLMs for more generic evolutionary optimization (Liu et al.,, 2023) but considers the case of conventional optimization and requires very detailed task information and descriptions. Closest to our approach is concurrently released work which also studies prompting strategies but targeting recombination operators for Evolution Strategies (ES) (Lange et al.,, 2024), where the LLM generates the next mean statistic of an ES which solutions are sampled from. Instead, our work considers the general QD problem which contains the added dimensions of diversity. Additionally, our approach taps into the curated solutions in the archive of elites to build the context and studies in-context prompt construction approaches for the QD problem and directly generates sample solutions.
@@
In QD, the goal is to find a diversity of high-quality solutions. This provides two dimensions of optimization, fitness and novelty/diversity, compared to conventional single-objective optimization approaches which only consider fitness. A common approach when using LLMs as optimizers is to sort the sequence in an increasing order of fitness and provide a query of higher value, allowing the model to complete the sequence (Yang et al.,, 2023; Mirchandani et al.,, 2023). However, in QD, we are not concerned with just finding a solution that has a high fitness, we are also interested in covering the feature space with high-scoring solutions. Figure 1 shows an overview of our approach to this, In-context QD. The LLM takes as input the current state of the archive, and uses this as a prompt to generate new solutions. The generated solutions are then considered for addition to grow the quality and diversity of the archive as in MAP-Elites. In the following, we explain how we represent the solutions and build a prompt that consists of the context and the query. This provides a general and modular framework for which to use LLMs as AI generators for Quality-Diversity.
@@
The solutions maintained in a QD archive lend themselves nicely to few-shot and many-shot prompting of LLMs as the diversity and quality of solutions provide a good set of examples and context to refer to in order to generate better ideas and solutions. Ideally, the entire archive could be used as the context. However, given the context size limit of many open-source LLMs, we use a sample/subset of solutions of the archive as an approximation of the archive to construct the context. If provided with models with larger context or problems with lower archive sizes, the entire archive could be used as the context. The context size is a hyper-parameter and can take values up to the archive size $C$. Yet, despite this temporary limitation, we can already demonstrate the effectiveness of this proposed approach.
In building the context, there are two considerations: the content and the structure of the context. As explained above, the content of the context is a subset representation of the archive. We uniformly sample solutions from the archive to form the content of the context. An increasingly larger context of solutions sampled would give us an increasingly better representation of the archive culminating in the entire archive itself. Meyerson et al., (2023) suggest that few-shot prompting can be viewed as Estimation of Distribution Algorithms (EDAs). In our QD setting, by providing more solutions from the archive in the context using our prompt template, we provide a good approximation of the patterns and the distributions of the fitness-feature space. This provides a better context for the language model especially when queried for desired features.
@@
The selection of features used in the query $d^{q}$ cannot be treated entirely the same as with the fitness. In the case of features, the goal is to cover the entire feature space and not find an optimal point in the space by moving in a specific direction of a feature. While expanding the horizon and boundary of the archive is needed, it is also just as important to fill in ”holes” in the archive to maximize the coverage of the feature space. An unbiased approach to this is to uniformly sample the centroids of the cells in the archive to obtain query features $d^{q}_{1},...,d^{q}_{G}$. We refer to this approach as ”Uniform”. We also propose a strategy which queries features from cells that are empty first, after which we default to sampling features uniformly from the cells in the archive when the number of empty cells becomes less than the batch size. The centroids of the empty cells and sampled cells are used for the queries. We refer to this strategy as ”Empty”.
@@
We compare in-context QD against a Random Sampling baseline which samples the parameter space uniformly and MAP-Elites (Mouret and Clune,, 2015) with the iso-line variation operator (Vassiliades and Mouret,, 2018). We evaluate the (i) maximum fitness score found throughout search, (ii) the coverage, measured by the number of filled cells in the archive and (iii) the QD score which is the sum of fitness of all the solutions in the archive. For all our algorithms and baselines, we use a batch size of $10$ at each generation, and run for 1000 generations. We also investigate the performance across different feature space sizes $C\in\{400,1600\}$ by changing the resolution of the MAP-Elites archive $C\in\{20^{2},40^{2}\}$ to ensure the generality of our approach. To be able to run in-context QD on just a single GPU device, we use Mistral-7B-v0.1 (Jiang et al.,, 2023), an open-source 7B parameter language model.